Shift report 20030125 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 25 2003 - 16:15:27 EST

  • Next message: Apache: "Shift report 20030125 16:00-24:00"
    ------------------------------------------------------------
    Date: Sat Jan 25 16:15:26 2003, Shift: 08:00-16:00
    Supervisor: PS, with: 
    Log book#13 pages 127 - 128
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam since 10:30, lost at 13:30 due to quench link interlock. Initial ZDC rate was 1700 at the end of run 900.
    
    2. DC scan runs: 7627 to 7633,
       run #7627 - HV_nom - 100V
       run #7628 - HV_nom - 50V
       run #7629 - HV_nom - 20V
       run #7630 - HV_nom 
       run #7631 - HV_nom +10(T3) + 20(T4/T5)
       run #7632 - HV_nom +20(T3) + 40(T4/T5)
       run #7633 - HV_nom +30(T3) + 60(T4/T5)
    
    Run 7634 is a normal production run. For all the data runs (7627 to 7633) D5 was off.
    
    Runs 7626, 7635 and 7636 are DVM calib., pedestal and TPC pedestal, respectively.
    
    After a preliminary analysis of DC scanning runs decided
    to keep HV=HV_nom+10(T3) +20(T4/T5).  
    
    Mapping change (affter run 7634) to removed the hot wire #48 in T3-1-V1.
    
    FV played with triggers see. elog for details.
    
    3. None detector problems. D5 tripped 
    twice in the last 24 hours. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Jan 25 2003 - 16:16:13 EST